# Fielding: Clustered Federated Learning with Data Drift

### v2

@@ 1 Introduction
As data drift accumulates, intra-cluster heterogeneity increases, and clusters can become as diverse as the full client set (Jothimurugesan et al., 2023; Moreno-Torres et al., 2012). Effective drift handling must address two key challenges: adapting to varying drift magnitudes while maintaining practical efficiency, and supporting diverse drift types.
@@
The second challenge is to detect and adapt to different types of drift. FL frameworks rely on client representations—client-side information used for clustering—to capture drifts. The label distribution vector is a common representation that reflects the label shift. Another is the input embedding: feature vectors derived from a client’s local data using a neural network, then aggregated (e.g., averaging) to represent the input distribution. When the input–output mapping remains stable, input feature shifts often correlate with label distribution shifts so that both representations can produce similar clustering results. However, neither captures changes in the input–output relationship itself – concept drift. Addressing concept drift typically requires loss-based representations, such as loss values, gradients, or gradient directions. For example, Sattler et al. (2021) clusters clients based on gradient directions to identify such shifts. Each representation has trade-offs: Label and embedding vectors capture only marginal distributions, while gradient-based features require additional computation, are sensitive to training noise, and evolve with model updates (Table 3 shows that gradient-based clustering generates better clusters as training progresses and the model becomes more stable). Selecting or combining these representations enables more reliable detection in all types of drift (see Appendix E).
@@ 2.1 Motivation and design choices
Per-client adjustment fails under many drift. Per-client adjustment methods, such as IFCA (Ghosh et al., 2020) and FlexCFL (Duan et al., 2021), maintain a fixed number of clusters and reassign drifted clients individually based on gradients or model weights. As mentioned in Section 1, such assignments become unstable and lead to poor clustering when many clients drift simultaneously. Figure 1 (“Move Individuals”) highlights this behavior. Between rounds 600–730, a significant drift event increases intra-cluster heterogeneity, and from rounds 740–1340, per-client adjustment fails to restore effective clustering. In fact, heterogeneity exceeds that of the unclustered client set.
@@ 2.3 Algorithmic framework
In this section, we describe our algorithm – presented in Algorithm 1 – in detail. Initially, Fielding clusters the clients using the $k$-means algorithm. The distance between clients is measured based on clients representations; in particular, for label-based clustering, we measure distance between data distributions, which is computed as the $\ell_{1}$-distance between their histograms: that is, assuming we have $l$ labels, the distance between clients with label distribution $p_{1},\ldots,p_{l}$ and $q_{1},\ldots,q_{l}$ is defined as $\sum_{i=1}^{l}|p_{i}-q_{i}|$.
We assume that the algorithm is executed for $T$ global iterations, and during each iteration, the label distribution of each client might change, for instance by adding or removing data points. The first step of Fielding is to handle data drift (Algorithm 2). Our goal is, by the end of each iteration, to maintain a good model for each cluster, with respect to its clients’ data. In this step, each drifted client is assigned to the closest cluster. If this causes a significant change in cluster centers – namely, if there exists a cluster moving by more than $\theta/3$, where $\theta$ is the average distance between cluster centers – we recluster clients from scratch. This condition avoids frequent reclustering, which might require excessive resources and can adversely change the clients’ losses. Importantly, we recluster all the clients, not just available clients, which, as we show in Appendix B, improves the accuracy.
@@ 2 Initialize cluster models $\mathbf{c}^{(1)},\ldots,\mathbf{c}^{(K)}$
4       Handle data drift using Algorithm 2
@@ 2.4 Convergence
In Appendix A, we analyze the performance of simplified version of our framework by bounding the average of clients’ loss functions at every iteration. The major challenge in the analysis is bounding the adverse effects of data drifts and reclustering. Data drifts change client datasets, hence changing their local objective functions. On the other hand, while reclustering is useful in the long run, its immediate effect on the objective can potentially be negative (see Figure 2(b)) due to the mixing of models from different clusters.
@@
Let $N$ be the number of clients and $M$ be the total number of machines sampled per round. Let $\mathbf{x}^{*}$ be the minimizer of $f_{0}=\operatorname*{avg}_{i}f^{(i)}_{0}$. Let $\mathbf{c}^{(k,*)}_{t}$ be the minimizer for cluster $k$ at iteration $t$. Then, under A, for $\eta\leq 1/L$, for any iteration $t$ we have
$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$ $\leq(1-\eta\mu)^{tR}\left(f_{0}(\mathbf{x}_{0})-f_{0}(\mathbf{x}^{*})\right)$
@@ 3 Evaluation
We evaluate Fielding prototype on four FL tasks with label distribution vectors as client representations (see Appendix E for results with other representations). End-to-end results highlight that Fielding improves final test accuracy by 1.9%-5.9% and is more stable than prior CFL methods. It works well with complementary FL optimizations, remains robust against malicious clients reporting corrupted distribution vectors, and still accelerates model convergence under low heterogeneity.
@@ 3.1 End-to-end training performance
Figure 4 shows that Fielding improves average client test accuracy by 5.9%, 1.9%, and 3.4% in FMoW, Cityscapes, and Waymo Open, respectively. These improvements are significant, given that baseline test accuracies range from 45% to 85%, and are comparable to the accuracy gains prior CFL works achieve on static data. When considering the global model’s final test accuracy as the target accuracy, Fielding offers a 1.27$\times$, 1.16$\times$, and 2.23$\times$ training speed-up, respectively (We define the point at which Fielding’s accuracy consistently surpasses the target accuracy as the moment it achieves the target accuracy). On the highly heterogeneous and dynamic Open Images trace, Fielding boosts accuracy by 17.9-26.1% in the final 100 rounds.
While Auxo performs comparably or slightly better than the global model baseline, the benefits are limited: up to 0.9% accuracy gain and up to 1.05$\times$ convergence speedup. These results fall short of those reported in the original paper on static datasets, suggesting a decline in clustering effectiveness. This degradation stems from re-clustering only the selected participants after each round, ignoring drifted but unselected clients. This approach also leads to sudden accuracy drops on the biased Waymo Open dataset, where over 80% samples are cars, due to sudden model shifts when divergent clients are selected. FlexCFL shows no improvement in accuracy or convergence across all real-world datasets, as it migrates clients individually without adjusting the number of clusters. As Figure 1 shows, this approach leads to cluster heterogeneity approaching that of the global client set.
FedDrift Comparison. FedDrift (Jothimurugesan et al., 2023) is another clustered FL method designed to handle drift. However, it is impractical for large-scale settings, as it (1) requires every client to train every cluster model in each round and (2) assumes all clients remain online and available for training. Therefore, we construct a small-scale setting with the FMoW dataset and 50 clients (the number of clients we selected per round in our original experiment). Figure 10 shows that Fielding achieves 2.5% higher final accuracy and reaches FedDrift’s final accuracy 2.01$\times$ faster.
@@ 3.2 Compatibility with other optimizations
We use the Oort selection algorithm (Lai et al., 2021) and a distance-based algorithm prioritizing clients closer to the distribution center as client selection examples. As shown in Figure 5, Fielding improves the final average test accuracy by 6.0% and 4.5% and reach the target accuracy 2.62×2.62$\times$2.62 × and 1.17×1.17$\times$1.17 × faster. Note that Oort selection actively incorporates clients otherwise filtered out due to long response time, leading to a longer average round time. Figure 6 shows that Fielding works well with aggregation algorithm FedYogi (Reddi et al., 2021) and q-FedAvg (Li et al., 2020c). Fielding improves the final accuracy by 5.9% and 4.9% while giving a 1.34×1.34$\times$1.34 × and 1.25×1.25$\times$1.25 × speedup respectively.
@@ 4 Related works
Clustered FL. FlexCFL (Duan et al., 2021) and IFCA (Ghosh et al., 2020) adjust clusters by relocating shifted clients while maintaining a fixed number of clusters, operating under the assumption of low drift magnitude. Auxo (Liu et al., 2023) and FedAC (Zhang et al., 2024) re-cluster selected clients using gradient information or local models, delaying drift adjustments for unselected clients. FedDrift (Jothimurugesan et al., 2023) re-clusters all clients by broadcasting all cluster models every round to compute client local training loss, leading to significant costs. Appendix B discusses other prior CFL works without drift handling and the impact of drifts on static clusters.
@@ Appendix A Proof of convergence
1 input: the number of clusters $K$, the number of data drift events $T$, the number of sampled clients per round $M$, the number of rounds per data drift event $R$, reclustering threshold $\Delta$
@@ 7             Let $\mathbf{x}_{i}$ be the model corresponding to the cluster containing client $i$
8      Let $r^{(1)}_{t},r^{(2)}_{t},\ldots$ be the client representations
@@ 21             $\mathbf{c}^{(k)}_{t+1}\leftarrow\tilde{\mathbf{c}}^{(k)}_{R}$
In this section, we analyze the performance of our framework in the case when the number of clusters $K$ is known and the number of local iterations is $1$. The analysis can be modified to handle local iterations and other variations based on the previous work, see e.g. Li et al. (2020a). To simplify the proof, in the algorithm, we make the following changes compared with the algorithm from the main body: 1) we use the $k$-center clustering instead of $k$-means clustering; 2) we perform reclustering when the intra-clsuter heterogeneity exceeds a certain threshold. We compare this threshold condition with the one from the main body in Section F.2.
We present the algorithm in Algorithm 4. After every data drift event, we first accept data changes for each client. After that, if the heterogeneity within some clsuter exceeds a certain threshold, we recluster the clients, and assign to each cluster a model by averaging models of all clients in the cluster. After that, for several rounds, we perform standard federated averaging updates on each cluster: we sample $M$ clients, compute the stochastic gradient for each client, and update the cluster model with the sampled gradients.
@@
where the last inequality follows by convexity of $f^{(i)}_{t}$. By F, since all clients $\bar{C}_{\ell}$ belong to the same cluster, their representations differ by at most $\Delta$, and hence by Assumption D their local objectives differ by at most $\theta\Delta$. Therefore,
@@
Hence, the sum of the first terms in Equation 1 over all $i$ can be bounded as
@@
Let $N$ be the number of clients, and let their objective functions satisfy Assumptions B-D. Let $M$ be the total number of machines sampled per round. Let $\mathbf{x}^{*}$ be the minimizer of $f_{0}=\operatorname*{avg}_{i}f^{(i)}_{0}$. Let $\mathbf{c}^{(k,*)}_{t}$ be the minimizer for cluster $k$ at data drift event $t$. Then, for $\eta\leq 1/L$, for any data drift event $t$ of Algorithm 4 we have
$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$ $\leq(1-\eta\mu)^{TR}\left(f(\mathbf{x}_{0})-f(\mathbf{x}^{*})\right)$
@@
$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$ $=\frac{1}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$
@@
Next, by Lemma 3, the total objective increases by at most $3\theta(\Delta+\delta)$ after reclustering. Since during $R$ rounds the non-stochastic part of each objective decreases by a factor of $(1-\eta\mu)^{R}$, we have
$\frac{1}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t+1}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t+1}(\mathbf{c}^{(k,*)}_{t+1})\right)$
@@
$\qquad+(1-\eta\mu)^{R}\frac{1}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$
By Lemma 2, the last (stochastic) term accumulates over all data drift events as $\frac{L\eta}{2\mu}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$. By a similar reasoning, $3\theta(\Delta+\delta)$ accumulates as $\frac{L\eta}{2\mu}\cdot 3\theta(\Delta+\delta)(1-\eta\mu)^{R}$. And finally, $f(\mathbf{x}_{0})-f(\mathbf{x}^{*})$ term is multiplied by $(1-\eta\mu)^{R}$ at every data drift event, giving factor of $(1-\eta\mu)^{TR}$ over $T$ data drift events. ∎
@@ B.2 Data drift affects static clustering
Data drift reduces the effectiveness of clustered FL as it increases data heterogeneity within clusters. We demonstrate this problem with the Functional Map of the World (FMoW) dataset (Christie et al., 2018). Recall from Section 3 that our preprocessed FmoW dataset contains time-stamped satellite images labeled taken on or after January 1, 2015, and each unique UTM zone metadata value corresponds to one client (302 clients in total). A day’s worth of data becomes available every two training rounds, and clients maintain images they received over the last 100 rounds.
As mentioned in Section 2, we use mean client distance as the intra-cluster heterogeneity. Note that instead of first averaging within each cluster and then finding the mean of those cluster averages, we use the overall mean across all clients. This helps us avoid biased results arising from imbalanced cluster sizes. Note that in the baseline case without any clustering, we consider all clients as members of a "global cluster".

### v3

@@ 1 Introduction
As data drift accumulates, intra‐cluster heterogeneity increases, and clusters can become as diverse as the full client set (Jothimurugesan et al., 2023; Moreno-Torres et al., 2012). Effective drift handling must address two key challenges: adapting to varying drift magnitudes while maintaining practical efficiency, and supporting diverse drift types.
@@
The second challenge is to detect and adapt to different types of drift. FL frameworks rely on client representations—client-side information used for clustering—to capture drifts. The label distribution vector is a common representation that reflects the label shift. Another is the input embedding: feature vectors derived from a client’s local data using a neural network, then aggregated (e.g., averaging) to represent the input distribution. When the input–output mapping remains stable, input feature shifts often correlate with label distribution shifts so that both representations can produce similar clustering results. However, neither captures changes in the input–output relationship itself – concept drift. Addressing concept drift typically requires loss-based representations, such as loss values, gradients, or gradient directions. For example, Sattler et al. (2021) clusters clients based on gradient directions to identify such shifts. Each representation has trade-offs: Label and embedding vectors capture only marginal distributions, while gradient-based features require additional computation, are sensitive to training noise, and evolve with model updates (Table 3 shows that gradient-based clustering generates better clusters as training progresses and the model becomes more stable). Selecting or combining these representations enables more reliable detection in all types of drift (see Appendix˜E).
@@ 2.1 Motivation and design choices
Per-client adjustment fails under many drift. Per-client adjustment methods, such as IFCA (Ghosh et al., 2020) and FlexCFL (Duan et al., 2021), maintain a fixed number of clusters and reassign drifted clients individually based on gradients or model weights. As mentioned in Section˜1, such assignments become unstable and lead to poor clustering when many clients drift simultaneously. Figure 1 (“Move Individuals”) highlights this behavior. Between rounds 600–730, a significant drift event increases intra-cluster heterogeneity, and from rounds 740–1340, per-client adjustment fails to restore effective clustering. In fact, heterogeneity exceeds that of the unclustered client set.
@@ 2.3 Algorithmic framework
In this section, we describe our algorithm – presented in Algorithm˜1 – in detail. Initially, Fielding clusters the clients using the $k$-means algorithm. The distance between clients is measured based on clients representations; in particular, for label-based clustering, we measure distance between data distributions, which is computed as the $\ell_{1}$-distance between their histograms: that is, assuming we have $l$ labels, the distance between clients with label distribution $p_{1},\ldots,p_{l}$ and $q_{1},\ldots,q_{l}$ is defined as $\sum_{i=1}^{l}|p_{i}-q_{i}|$.
We assume that the algorithm is executed for $T$ global iterations, and during each iteration, the label distribution of each client might change, for instance by adding or removing data points. The first step of Fielding is to handle data drift (Algorithm˜2). Our goal is, by the end of each iteration, to maintain a good model for each cluster, with respect to its clients’ data. In this step, each drifted client is assigned to the closest cluster. If this causes a significant change in cluster centers – namely, if there exists a cluster moving by more than $\theta/3$, where $\theta$ is the average distance between cluster centers – we recluster clients from scratch. This condition avoids frequent reclustering, which might require excessive resources and can adversely change the clients’ losses. Importantly, we recluster all the clients, not just available clients, which, as we show in Appendix˜B, improves the accuracy.
@@ 2 Initialize cluster models $\mathbf{c}^{(1)},\ldots,\mathbf{c}^{(K)}$
4   Handle data drift using Algorithm˜2
@@ 2.4 Convergence
In Appendix˜A, we analyze the performance of simplified version of our framework by bounding the average of clients’ loss functions at every iteration. The major challenge in the analysis is bounding the adverse effects of data drifts and reclustering. Data drifts change client datasets, hence changing their local objective functions. On the other hand, while reclustering is useful in the long run, its immediate effect on the objective can potentially be negative (see Figure˜2(b)) due to the mixing of models from different clusters.
@@
Let $N$ be the number of clients and $M$ be the total number of machines sampled per round. Let $\mathbf{x}^{*}$ be the minimizer of $f_{0}=\operatorname*{avg}_{i}f^{(i)}_{0}$. Let $\mathbf{c}^{(k,*)}_{t}$ be the minimizer for cluster $k$ at iteration $t$. Then, under ˜A, for $\eta\leq 1/L$, for any iteration $t$ we have
[…]
@@ 3 Evaluation
We evaluate Fielding prototype on four FL tasks with label distribution vectors as client representations (see Appendix˜E for results with other representations). End-to-end results highlight that Fielding improves final test accuracy by 1.9%-5.9% and is more stable than prior CFL methods. It works well with complementary FL optimizations, remains robust against malicious clients reporting corrupted distribution vectors, and still accelerates model convergence under low heterogeneity.
@@ 3.1 End-to-end training performance
[…]
@@ 3.2 Compatibility with other optimizations
We use the Oort selection algorithm (Lai et al., 2021) and a distance-based algorithm prioritizing clients closer to the distribution center as client selection examples. As shown in Figure 5, Fielding improves the final average test accuracy by 6.0% and 4.5% and reach the target accuracy $2.62\times$ and $1.17\times$ faster. Note that Oort selection actively incorporates clients otherwise filtered out due to long response time, leading to a longer average round time. Figure 6 shows that Fielding works well with aggregation algorithm FedYogi (Reddi et al., 2021) and q-FedAvg (Li et al., 2020c). Fielding improves the final accuracy by 5.9% and 4.9% while giving a $1.34\times$ and $1.25\times$ speedup respectively.
@@ 4 Related works
Clustered FL. FlexCFL (Duan et al., 2021) and IFCA (Ghosh et al., 2020) adjust clusters by relocating shifted clients while maintaining a fixed number of clusters, operating under the assumption of low drift magnitude. Auxo (Liu et al., 2023) and FedAC (Zhang et al., 2024) re-cluster selected clients using gradient information or local models, delaying drift adjustments for unselected clients. FedDrift (Jothimurugesan et al., 2023) re-clusters all clients by broadcasting all cluster models every round to compute client local training loss, leading to significant costs. Appendix˜B discusses other prior CFL works without drift handling and the impact of drifts on static clusters.
@@ Appendix A Proof of convergence
1input: the number of clusters $K$, the number of data drift events $T$, the number of sampled clients per round $M$, the number of rounds per data drift event $R$, reclustering threshold $\Delta$
@@ 7      Let $\mathbf{x}_{i}$ be the model corresponding to the cluster containing client $i$
8  Let $r^{(1)}_{t},r^{(2)}_{t}$$\ldots$ be the client representations
@@ 21    $\mathbf{c}^{(k)}_{t+1}\leftarrow\tilde{\mathbf{c}}^{(k)}_{R}$
In this section, we analyze the performance of our framework in the case when the number of clusters $K$ is known and the number of local iterations is $1$. The analysis can be modified to handle local iterations and other variations based on the previous work, see e.g. Li et al. (2020a). To simplify the proof, in the algorithm, we make the following changes compared with the algorithm from the main body: 1) we use the $k$-center clustering instead of $k$-means clustering; 2) we perform reclustering when the intra-clsuter heterogeneity exceeds a certain threshold. We compare this threshold condition with the one from the main body in Section˜F.2.
We present the algorithm in Algorithm˜4. After every data drift event, we first accept data changes for each client. After that, if the heterogeneity within some clsuter exceeds a certain threshold, we recluster the clients, and assign to each cluster a model by averaging models of all clients in the cluster. After that, for several rounds, we perform standard federated averaging updates on each cluster: we sample $M$ clients, compute the stochastic gradient for each client, and update the cluster model with the sampled gradients.
@@
where the last inequality follows by convexity of $f^{(i)}_{t}$. By ˜F, since all clients $\bar{C}_{\ell}$ belong to the same cluster, their representations differ by at most $\Delta$, and hence by Assumption D their local objectives differ by at most $\theta\Delta$. Therefore,
@@
Hence, the sum of the first terms in Equation˜1 over all $i$ can be bounded as
@@
Let $N$ be the number of clients, and let their objective functions satisfy Assumptions B-D. Let $M$ be the total number of machines sampled per round. Let $\mathbf{x}^{*}$ be the minimizer of $f_{0}=\operatorname*{avg}_{i}f^{(i)}_{0}$. Let $\mathbf{c}^{(k,*)}_{t}$ be the minimizer for cluster $k$ at data drift event $t$. Then, for $\eta\leq 1/L$, for any data drift event $t$ of Algorithm˜4 we have
[…]
@@
[…]
@@
Next, by Lemma˜3, the total objective increases by at most $3\theta(\Delta+\delta)$ after reclustering. Since during $R$ rounds the non-stochastic part of each objective decreases by a factor of $(1-\eta\mu)^{R}$, we have
[…]
@@
[…]
By Lemma˜2, the last (stochastic) term accumulates over all data drift events as $\frac{L\eta}{2\mu}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$. By a similar reasoning, $3\theta(\Delta+\delta)$ accumulates as $\frac{L\eta}{2\mu}\cdot 3\theta(\Delta+\delta)(1-\eta\mu)^{R}$. And finally, $f(\mathbf{x}_{0})-f(\mathbf{x}^{*})$ term is multiplied by $(1-\eta\mu)^{R}$ at every data drift event, giving factor of $(1-\eta\mu)^{TR}$ over $T$ data drift events. ∎
@@ B.2 Data drift affects static clustering
Data drift reduces the effectiveness of clustered FL as it increases data heterogeneity within clusters. We demonstrate this problem with the Functional Map of the World (FMoW) dataset (Christie et al., 2018). Recall from Section˜3 that our preprocessed FmoW dataset contains time-stamped satellite images labeled taken on or after January 1, 2015, and each unique UTM zone metadata value corresponds to one client (302 clients in total). A day’s worth of data becomes available every two training rounds, and clients maintain images they received over the last 100 rounds.
As mentioned in Section˜2, we use mean client distance as the intra-cluster heterogeneity. Note that instead of first averaging within each cluster and then finding the mean of those cluster averages, we use the overall mean across all clients. This helps us avoid biased results arising from imbalanced cluster sizes. Note that in the baseline case without any clustering, we consider all clients as members of a "global cluster".